OpenStudy (anonymous):

Help me,por favor? The problem is (a^2b^3)(ab)^-2 and the answer is b,but how do I get there? step by step answer,plz and thank you!

OpenStudy (anonymous):

\[(a^2b^3)(ab)^-2= \frac{a^2b^3}{(ab)^2}\]

OpenStudy (anonymous):

Note: \[x^{-n}= \frac{1}{x^n}\]

OpenStudy (anonymous):

\[(a^2b^3)(ab)^{-2}= \frac{a^2b^3}{(ab)^2}\] \[= b\]

OpenStudy (anonymous):

why is it b though???

OpenStudy (anonymous):

\[(a^2b^3)(ab)^{-2}= \frac{a^2b^3}{(ab)^2}= \frac{a^2b^3}{a^2b^2}=b\]

OpenStudy (anonymous):

\[Note:(ab)^2=a^2b^2\]

OpenStudy (anonymous):

OH!!!! Okay,thanks!! Your fantabulastic :)

OpenStudy (anonymous):

\[\frac{a^2}{a^2}=1 \ \ \frac{b^3}{b^2}=b \]

OpenStudy (anonymous):

You are Welcome :)

OpenStudy (anonymous):

Wait 1 sec,DONT LEAVE.Will post 1 more question,but a new 1 so I can give you another medal

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

If you want you can post it here itself i will help